OpenStudy (anonymous):

If y varies directly as the cube of x, what is the value of y in these ordered pairs? (4,16) and (2, y)

OpenStudy (johnweldon1993):

Well if y varies directly as the cube of x we have \[y = kx^3\] We know that in that first ordered pair x = 4 and y = 16 so \[16 = k\times4^3\] A little simplification and we get \[k = \frac{ 16 }{ 64 }\] More simplified down we get \[k = \frac{ 1 }{ 4 }\] So now that we have what 'k' equals...we can figure out what 'y' equals in that 2nd ordered pair \[y = kx^3\] plug in everything we know \[y = \frac{ 1 }{ 4 }\times2^3\] What does 'y' equal...?

OpenStudy (anonymous):

Y=2?

OpenStudy (johnweldon1993):

That would be correct!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (johnweldon1993):

No problem!